# Semantic Mapping in Indoor Embodied AI – A Survey on Advances, Challenges, and Future Directions

### v2

@@ 4.3 Dense geometric map
Semantic information can also be directly accumulated onto the 3D geometry of a scene using triangle meshes (Valentin et al., 2013), surfels (Stückler & Behnke, 2014) or point clouds (Jatavallabhula et al., 2023), with point cloud maps being the most widely used due to simplicity and ease of use. While these representations are not always traditionally considered “maps”, they effectively function as semantic maps by coupling geometric structure with semantic content. Dense geometric map representations, particularly point clouds, have long been widely used in robotics for tasks such as mapping, localization and navigation, as well as in broader 3D scene understanding tasks (Peng et al., 2023; Xu et al., 2024). Recently, there has been a surge of interest in applying these representations to embodied AI for downstream reasoning and decision-making (Gu et al., 2023b). We further categorize dense geometric map representations into point cloud maps (Sec. 4.3.1) and neural fields (Sec. 4.3.2). Point cloud maps store 3D points with associated semantic labels whereas neural fields represent scenes as continuous functions using neural networks (see Tab. 5 for a summary of prior works).
@@ 4.3.1 Point cloud maps
In robotics, point cloud-based mapping has undergone significant evolution, driven by advances in sensing, computational power and SLAM algorithms. Early efforts (Nüchter et al., 2007; May et al., 2009) use 3D sensor scans to build geometrically consistent representations of large-scale environments by emphasizing accuracy and loop-closure. As real-time visual SLAM has matured, dense point cloud generation from RGB-D sensors became feasible, enabling detailed indoor reconstructions (Whelan et al., 2015; Dai et al., 2017; Schops et al., 2019; Min & Dunn, 2021). More recently, learning-based SLAM systems like Droid-SLAM (Teed & Deng, 2021) integrate deep feature extraction and end-to-end optimization to improve SLAM. CoFusion (Rünz & Agapito, 2017) introduce systems for dense SLAM with semantic fusion, pushing toward maps that are not only geometrically accurate but also semantically meaningful. This trend has been accelerated recently with ConceptFusion (Jatavallabhula et al., 2023) and Clio (Maggio et al., 2024), which align open-vocabulary visual features (e.g., CLIP) with 3D reconstructions to create task-agnostic and queryable maps. These efforts represent a shift from purely geometric maps toward dense, semantically-rich, open-vocabulary representations that support a wide range of downstream tasks. Moreover, Gaussian splatting has recently emerged as a powerful alternative to traditional point cloud representations in robotics, enabling high-fidelity scene reconstruction and differentiable 3D representations useful for tasks such as visual localization, SLAM, and manipulation. Recent works such as SGS-SLAM (Li et al., 2024a) demonstrate how to include semantic labels onto Gaussian splats to effectively perform SLAM, and GaussianGrasper (Zheng et al., 2024b) demonstrates how these representations can be fused with semantics to support open-vocabulary grasping.
@@ 4.4 Hybrid maps
The combination of metric information from grid-based maps together with a topological map is also known as a topometric map (Thrun et al., 1998b; Tomatis et al., 2001; Blanco et al., 2008b; Konolige et al., 2011; Ko et al., 2013; An et al., 2023) (Mozos et al., 2007; Vasudevan & Siegwart, 2008; Zender et al., 2008; Pronobis et al., 2010; Hemachandra et al., 2011; Walter et al., 2013; 2014; Hemachandra et al., 2015; Duvallet et al., 2016; Walter et al., 2022). Thrun et al. (1998b) proposes a single statistical mapping algorithm that first constructs a coarse topological map and uses it to construct a fine-grained grid map. They show that the topological map solves a global alignment problem by correcting large odometry errors, while the grid map solves a local alignment problem by producing high-resolution maps. Tomatis et al. (2001), on the other hand, proposes a compact environmental model where corners and hallways are represented by a topological map and rooms are represented by a grid map, both of which are connected in a single representation. When the robot is moving in hallways, it creates and updates the global topological map, and as soon as it enters a room, it creates a new local metric map [^1] They differentiate hallways and rooms by using laser sensor such that thin long open spaces are considered as hallways whereas other open spaces are considered as rooms.. They argue that the robot will only need to be precise inside rooms (e.g. manipulating objects, etc.) which justifies the need for the fine-grained precise metric maps, while the topological map is used to simply maintain global consistency in indistinguishable spaces such as long hallways and transitioning between significant places. BEVBert (An et al., 2023) is a more recent method that constructs hybrid maps offline and then learn a multimodal map representation to perform better spatial reasoning in the complex task of language-guided navigation.
@@ 5 Map encoding
In this section we will discuss different ways information is encoded and stored into semantic maps. Irrespective of how the map is structured, the map encoding, i.e. the values stored in the map can be either explicit or implicit. An explicit map encoding is one where the type of information stored is clearly known. An implicit encoding, on the other hand, uses a feature embedding that may not be directly interpretable. We now summarize various works that explore these two types of encodings (Fig. 7).
@@ 5.1 Explicit encoding
Many prior works store explicit information in the map, such as occupancy, which has proven beneficial in obstacle avoidance (Elfes, 1989; May et al., 2009; Chaplot et al., 2019; Georgakis et al., 2022a). In such cases, the spatial map stores a binary occupancy value of $1$ or 00 depending on whether the corresponding location in the physical environment is ‘occupied’ by objects or not.
@@
While the above works build explicit spatial maps, some works also build explicit topological maps (Choset & Nagatani, 2001; Blöchliger et al., 2018; Chen et al., 2022; Salas-Moreno et al., 2013; Duvallet et al., 2013; Patki et al., 2019; Arkin et al., 2020a; Raychaudhuri et al., 2025; Chang et al., 2025). Kwon et al. (2021) store the visitation timestep in the graph node in order to encode a temporal relation between visited locations. This information is then replaced by the latest visitation timestep while aggregating the map.
@@ 5.2.1 Closed-vocabulary encoding
ConceptGraphs (2023b) AI2Thor, robot CLIP, DINO GPT-4 highest similarity score
@@ 5.2.2 Open-vocabulary encoding
The limitation of the closed-vocabulary encoding can be mitigated by extracting features from a Large Vision-Language Model (LVLM) that was jointly trained on a vast amount of internet data of images and their text captions, such as CLIP (Radford et al., 2021). This allows the map to store information about ‘any’ object in the environment and eventually be queried via an open-vocabulary text query (Wen et al., 2025) not limited to a predefined set of object categories. For example, an agent may be asked to ‘find a red and blue striped zebra toy in the children’s room’. It is likely that the agent has not seen a ‘red and blue striped zebra toy’ during training but can leverage a LVLM to reason about its prior knowledge about zebras, colors and rooms in general. Moreover, recently large language models (LLMs), such as GPT-4 (OpenAI, 2023), have been shown to be able to perform complex task planning. Thus open-vocabulary maps built using LVLM along with LLM-based planners (Taioli et al., 2023; Huang et al., 2023b; Long et al., 2024) have led to a recent line of works in embodied agents (Tab. 6, Fig. 8).
@@
Figure 8: Open-Vocabulary map building. There has been a growing interest to build flexible open-vocabulary maps which can be built once and then used in various downstream tasks during inference. VLMap (Huang et al., 2023a) and NLMap (Chen et al., 2023a) structure their maps as spatial maps while ConceptGraphs (Gu et al., 2023b) build a topological map. (Figures reproduced from paper.)
To mitigate this issue, researchers need to develop methods that identify where objects are located in the image, and then extract features for those objects. One way to achieve this is to store feature embeddings for each pixel in the input image. This can be done using LSeg (Li et al., 2022) which outputs pixel-level embeddings given an image as is done in VLMaps (Huang et al., 2023a). The embeddings are then projected into a 2D spatial map using depth observations. When multiple points are projected onto the same grid cell on the map, VLMaps aggregates by averaging the features. During inference, they extract object names from the language query and calculate pixel-text similarity scores on the map to retrieve the objects of interest. However, these pixel-level embeddings are very dense and can be redundant since not all pixels in an image contain objects. Second, some information might be lost while averaging the pixel-level embeddings. A third issue is that it does not encode object-level semantics, thus ignoring information about spatial relationships. OneMap (Busch et al., 2024) stores patch-level features extracted from SED (Xie et al., 2024) with a hierarchical encoder-based backbone which has shown to capture spatial information better than the transformer-based architecture. This alleviates the issues of using pixel-based features.
@@
While NLMap builds a 3D spatial map, many (Shah et al., 2023; Garg et al., 2024) build a topological map following a similar approach. For instance, ConceptGraphs (Gu et al., 2023b) is a recent approach that retrieves the objects of interest from the input image by using the class-agnostic 2D segmentation model, Segment-Anything (SAM) (Kirillov et al., 2023) to obtain candidate masks. These objects then form the nodes of the topological map. The image features for each object can then be extracted by CLIP and DINO (Oquab et al., 2023) to be stored in the corresponding node. Additional information about the objects can also be stored in each node. For example, ConceptGraphs store the point cloud of the proposed masks and a caption for each object as obtained by using LLaVA (Liu et al., 2023a) and GPT-4 (OpenAI, 2023) along with a point cloud obtained by projecting the object mask proposed by SAM into 3D space. After the nodes are formed, edges between the nodes can be constructed depending on whether the objects are spatially related. In ConceptGraphs, spatial relation between two objects is determined by whether the point clouds of the respective objects have a geometric similarity or overlap. In other words, when a certain proportion of points in point cloud of one object lie within a distance threshold of that of the second object, the objects can be said to be spatially related to each other and an edge is constructed between the corresponding nodes. The edges additionally store a spatial relationship description obtained from LLM. When a newly detected object is found to be similar to a node, it is updated with averaged object features, union of point clouds and the latest object caption. In case it is not similar, a new node is added to the graph. ConceptGraphs show that a single map representation built in this manner can be successfully and effectively used in object grounding, robot navigation and robot manipulation tasks.
@@ 6.1 Extrinsic evaluation
This section discusses extrinsic (or task-level) evaluation metrics commonly used to assess agent performance across embodied tasks. In navigation, success is not simply reaching the goal but signaling it intentionally, often via a dedicated action like ‘Stop’ or ‘Found’ (Anderson et al., 2018a). Accordingly, Success Rate (SR) is used to evaluate whether the agent correctly signals goal completion, while Success weighted by inverse Path Length (SPL) accounts for trajectory efficiency. These metrics are standard in navigation benchmarks (Beeching et al., 2020; Batra et al., 2020b; Chaplot et al., 2020a; b; Wijmans et al., 2019b; Li et al., 2021c; Krantz et al., 2022; Georgakis et al., 2022a; Taioli et al., 2024; Yokoyama et al., 2024). Additional metrics include Oracle SR (OSR), which accounts for when the agent reaches the goal but fails to signal, and Navigation Error (NE), which measures the final distance to the goal. In multi-step tasks or multi-goal navigation, metrics like Progress (Wani et al., 2020; Khanna et al., 2024) or Goal-Condition Success (Shridhar et al., 2020; Padmakumar et al., 2022) capture the proportion of successfully completed subgoals. For instruction following, normalized Dynamic Time Warping (nDTW) is used to evaluate how well the agent’s trajectory aligns with the reference path (Ilharco et al., 2019; Krantz et al., 2020; Raychaudhuri et al., 2021; 2025). Tasks involving classification or detection often use precision (how often predictions are correct), recall (how well all relevant targets are found), F1 score (combines and balances precision and recall) and receiver-operating characteristic curve or ROC curve (measures the tradeoff between sensitivity or true positive rate and specificity or false positive rate) (Chen & Mooney, 2011a; Magnusson et al., 2009; Matuszek et al., 2012; Tellex et al., 2011b; Wellhausen et al., 2020) to evaluate prediction quality. In exploration tasks, coverage is used to measure the proportion of the environment observed (Chaplot et al., 2019). In manipulation tasks such as object rearrangement or tabletop operations, task success rate, measuring the percentage of trials where the agent achieves the desired manipulation (e.g., grasping, placing), is the dominant metric (Kalashnikov et al., 2018; Batra et al., 2020a; Lin et al., 2020; Zeng et al., 2020; Huang et al., 2023b; Gu et al., 2023a). Some works also evaluate policy outputs using precision (Zeng et al., 2017), or employ pose error metrics to assess deviations between the final and target object positions (Liu et al., 2012).
@@ 6.2 Intrinsic evaluation
Accuracy. Map accuracy refers to how accurately the map captures geometric or semantic information when compared against the ground truth. However obtaining the ground truth map can be challenging in most cases. Some works (Georgakis et al., 2022a; Ramakrishnan et al., 2020) evaluate the 2D occupancy maps by using map accuracy to measure how much of that area matches with the ground-truth, and IoU to measure intersection over union between the built and the ground-truth map. While others (Cartillier et al., 2021; Georgakis et al., 2022b) evaluate the accuracy of the built 2D semantic map using semantic segmentation metrics such as pixel-wise labeling accuracy, pixel-based F1 score, IoU score and a contour-based average boundary F1 score. However, there is a lack of standardized metrics that can be used across various map types. For example, IoU is not directly applicable to topological maps. Moreover, while evaluating accuracy for semantic maps that store a fixed set of object categories is straightforward, it is tricky to evaluate semantic maps that store open-vocabulary features. As an alternative, OpenScene (Peng et al., 2023) performs semantic segmentation on their open-vocabulary map for a fixed set of categories and thereafter evaluate it using accuracy and IoU. ConceptGraphs (Gu et al., 2023b), on the other hand, evaluate their open-vocabulary topological map by employing human evaluators on Amazon Mechanical Turk (AMT) to report the scene graph accuracy. OpenLex3D (Kassab et al., 2025) has been recently introduced to evaluate open-vocabulary 3D scene representations, by providing object labels at various granularity or precision. Similar evaluation techniques could benefit intrinsic evaluation in embodied setting as well. The challenge to evaluate modern-day open-vocabulary maps lies in collecting precise ground-truth maps for large real-world environments, which is labor-intensive or infeasible. This limits the use of direct comparisons in order to calculate map accuracy. Moreover, accuracy in dynamic or cluttered environments is also rarely assessed. Both these provide possible future area of research.
@@
Precision, Recall, F1, ROC Language grounding (2011a; 2012; 2011b)

### v3

@@ 4.3 Dense geometric map
Semantic information can also be directly accumulated onto the 3D geometry of a scene using triangle meshes (Valentin et al., 2013), surfels (Stückler & Behnke, 2014) or point clouds (Jatavallabhula et al., 2023), with point cloud maps being the most widely used due to simplicity and ease of use. While these representations are not always traditionally considered “maps”, they effectively function as semantic maps by coupling geometric structure with semantic content. Dense geometric map representations, particularly point clouds, have long been widely used in robotics for tasks such as mapping, localization and navigation, as well as in broader 3D scene understanding tasks (Peng et al., 2023; Xu et al., 2024). Recently, there has been a surge of interest in applying these representations to embodied AI for downstream reasoning and decision-making (Gu et al., 2024). We further categorize dense geometric map representations into point cloud maps (Sec.˜4.3.1) and neural fields (Sec.˜4.3.2). Point cloud maps store 3D points with associated semantic labels whereas neural fields represent scenes as continuous functions using neural networks (see Tab.˜5 for a summary of prior works).
@@ 4.3.1 Point cloud maps
In robotics, point cloud-based mapping has undergone significant evolution, driven by advances in sensing, computational power and SLAM algorithms. Early efforts (Nüchter et al., 2007; May et al., 2009) use 3D sensor scans to build geometrically consistent representations of large-scale environments by emphasizing accuracy and loop-closure. As real-time visual SLAM has matured, dense point cloud generation from RGB-D sensors became feasible, enabling detailed indoor reconstructions (Whelan et al., 2015; Dai et al., 2017; Schops et al., 2019; Min & Dunn, 2021). More recently, learning-based SLAM systems like Droid-SLAM (Teed & Deng, 2021) integrate deep feature extraction and end-to-end optimization to improve SLAM. CoFusion (Rünz & Agapito, 2017) introduce systems for dense SLAM with semantic fusion, pushing toward maps that are not only geometrically accurate but also semantically meaningful. This trend has been accelerated recently with ConceptFusion (Jatavallabhula et al., 2023), which aligns open-vocabulary visual features (e.g., CLIP) with 3D reconstructions to create task-agnostic and queryable maps. These efforts represent a shift from purely geometric maps toward dense, semantically-rich, open-vocabulary representations that support a wide range of downstream tasks. Moreover, Gaussian splatting has recently emerged as a powerful alternative to traditional point cloud representations in robotics, enabling high-fidelity scene reconstruction and differentiable 3D representations useful for tasks such as visual localization, SLAM, and manipulation. Recent works such as SGS-SLAM (Li et al., 2024a) demonstrate how to include semantic labels onto Gaussian splats to effectively perform SLAM, and GaussianGrasper (Zheng et al., 2024b) demonstrates how these representations can be fused with semantics to support open-vocabulary grasping.
@@ 4.4 Hybrid maps
The combination of metric information from grid-based maps together with a topological map is also known as a topometric map (Thrun et al., 1998b; Tomatis et al., 2001; Blanco et al., 2008b; Konolige et al., 2011; Ko et al., 2013; An et al., 2023; Mozos et al., 2007; Vasudevan & Siegwart, 2008; Zender et al., 2008; Pronobis et al., 2010; Hemachandra et al., 2011; Walter et al., 2013; 2014; Hemachandra et al., 2015; Duvallet et al., 2016; Walter et al., 2022). Thrun et al. (1998b) proposes a single statistical mapping algorithm that first constructs a coarse topological map and uses it to construct a fine-grained grid map. They show that the topological map solves a global alignment problem by correcting large odometry errors, while the grid map solves a local alignment problem by producing high-resolution maps. Tomatis et al. (2001), on the other hand, proposes a compact environmental model where corners and hallways are represented by a topological map and rooms are represented by a grid map, both of which are connected in a single representation. When the robot is moving in hallways, it creates and updates the global topological map, and as soon as it enters a room, it creates a new local metric map [^1] They differentiate hallways and rooms by using laser sensor such that thin long open spaces are considered as hallways whereas other open spaces are considered as rooms.. They argue that the robot will only need to be precise inside rooms (e.g. manipulating objects, etc.) which justifies the need for the fine-grained precise metric maps, while the topological map is used to simply maintain global consistency in indistinguishable spaces such as long hallways and transitioning between significant places. BEVBert (An et al., 2023) is a more recent method that constructs hybrid maps offline and then learn a multimodal map representation to perform better spatial reasoning in the complex task of language-guided navigation.
@@ 5 Map encoding
In this section we will discuss different ways information is encoded and stored into semantic maps. Irrespective of how the map is structured, the map encoding, i.e. the values stored in the map can be either explicit or implicit. An explicit map encoding is one where the type of information stored is clearly known. An implicit encoding, on the other hand, uses a feature embedding that may not be directly interpretable. We now summarize various works that explore these two types of encodings (Fig.˜8).
@@ 5.1 Explicit encoding
Many prior works store explicit information in the map, such as occupancy, which has proven beneficial in obstacle avoidance (Elfes, 1989; May et al., 2009; Chaplot et al., 2019; Georgakis et al., 2022a). In such cases, the spatial map stores a binary occupancy value of $1$ or 0 depending on whether the corresponding location in the physical environment is ‘occupied’ by objects or not.
@@
While the above works build explicit spatial maps, some works also build explicit topological maps (Choset & Nagatani, 2001; Blochliger et al., 2018; Chen et al., 2022; Salas-Moreno et al., 2013; Duvallet et al., 2013; Patki et al., 2019; Arkin et al., 2020a; Raychaudhuri et al., 2025; Chang et al., 2025). Kwon et al. (2021) store the visitation timestep in the graph node in order to encode a temporal relation between visited locations. This information is then replaced by the latest visitation timestep while aggregating the map.
@@ 5.2.1 Closed-vocabulary encoding
ConceptGraphs (2024) AI2Thor, robot CLIP, DINO GPT-4 highest similarity score
@@ 5.2.2 Open-vocabulary encoding
The limitation of the closed-vocabulary encoding can be mitigated by extracting features from a Large Vision-Language Model (LVLM) that was jointly trained on a vast amount of internet data of images and their text captions, such as CLIP (Radford et al., 2021). This allows the map to store information about ‘any’ object in the environment and eventually be queried via an open-vocabulary text query (Wen et al., 2025) not limited to a predefined set of object categories. For example, an agent may be asked to ‘find a red and blue striped zebra toy in the children’s room’. It is likely that the agent has not seen a ‘red and blue striped zebra toy’ during training but can leverage a LVLM to reason about its prior knowledge about zebras, colors and rooms in general. Moreover, recently large language models (LLMs), such as GPT-4 (OpenAI, 2023), have been shown to be able to perform complex task planning. Thus open-vocabulary maps built using LVLM along with LLM-based planners (Taioli et al., 2023; Huang et al., 2023b; Long et al., 2024) have led to a recent line of works in embodied agents (Tab.˜6, Fig.˜9).
@@
Figure 9: Open-Vocabulary map building. There has been a growing interest to build flexible open-vocabulary maps which can be built once and then used in various downstream tasks during inference. VLMap (Huang et al., 2023a) and NLMap (Chen et al., 2023a) structure their maps as spatial maps while ConceptGraphs (Gu et al., 2024) build a topological map. (Figures reproduced from paper.)
To mitigate this issue, researchers need to develop methods that identify where objects are located in the image, and then extract features for those objects. One way to achieve this is to store feature embeddings for each pixel in the input image. This can be done using LSeg (Li et al., 2022) which outputs pixel-level embeddings given an image as is done in VLMaps (Huang et al., 2023a). The embeddings are then projected into a 2D spatial map using depth observations. When multiple points are projected onto the same grid cell on the map, VLMaps aggregates by averaging the features. During inference, they extract object names from the language query and calculate pixel-text similarity scores on the map to retrieve the objects of interest. However, these pixel-level embeddings are very dense and can be redundant since not all pixels in an image contain objects. Second, some information might be lost while averaging the pixel-level embeddings. A third issue is that it does not encode object-level semantics, thus ignoring information about spatial relationships. OneMap (Busch et al., 2025) stores patch-level features extracted from SED (Xie et al., 2024) with a hierarchical encoder-based backbone which has shown to capture spatial information better than the transformer-based architecture. This alleviates the issues of using pixel-based features.
@@
While NLMap builds a 3D spatial map, many (Shah et al., 2023; Garg et al., 2024) build a topological map following a similar approach. For instance, ConceptGraphs (Gu et al., 2024) is a recent approach that retrieves the objects of interest from the input image by using the class-agnostic 2D segmentation model, Segment-Anything (SAM) (Kirillov et al., 2023) to obtain candidate masks. These objects then form the nodes of the topological map. The image features for each object can then be extracted by CLIP and DINO (Oquab et al., 2023) to be stored in the corresponding node. Additional information about the objects can also be stored in each node. For example, ConceptGraphs store the point cloud of the proposed masks and a caption for each object as obtained by using LLaVA (Liu et al., 2023a) and GPT-4 (OpenAI, 2023) along with a point cloud obtained by projecting the object mask proposed by SAM into 3D space. After the nodes are formed, edges between the nodes can be constructed depending on whether the objects are spatially related. In ConceptGraphs, spatial relation between two objects is determined by whether the point clouds of the respective objects have a geometric similarity or overlap. In other words, when a certain proportion of points in point cloud of one object lie within a distance threshold of that of the second object, the objects can be said to be spatially related to each other and an edge is constructed between the corresponding nodes. The edges additionally store a spatial relationship description obtained from LLM. When a newly detected object is found to be similar to a node, it is updated with averaged object features, union of point clouds and the latest object caption. In case it is not similar, a new node is added to the graph. ConceptGraphs show that a single map representation built in this manner can be successfully and effectively used in object grounding, robot navigation and robot manipulation tasks.
@@ 6.1 Extrinsic evaluation
This section discusses extrinsic (or task-level) evaluation metrics commonly used to assess agent performance across embodied tasks. In navigation, success is not simply reaching the goal but signaling it intentionally, often via a dedicated action like ‘Stop’ or ‘Found’ (Anderson et al., 2018a). Accordingly, Success Rate (SR) is used to evaluate whether the agent correctly signals goal completion, while Success weighted by inverse Path Length (SPL) accounts for trajectory efficiency. These metrics are standard in navigation benchmarks (Beeching et al., 2020; Batra et al., 2020b; Chaplot et al., 2020a; b; Wijmans et al., 2019b; Li et al., 2021c; Krantz et al., 2022; Georgakis et al., 2022a; Taioli et al., 2024; Yokoyama et al., 2024). Additional metrics include Oracle SR (OSR), which accounts for when the agent reaches the goal but fails to signal, and Navigation Error (NE), which measures the final distance to the goal. In multi-step tasks or multi-goal navigation, metrics like Progress (Wani et al., 2020; Khanna et al., 2024) or Goal-Condition Success (Shridhar et al., 2020; Padmakumar et al., 2022) capture the proportion of successfully completed subgoals. For instruction following, normalized Dynamic Time Warping (nDTW) is used to evaluate how well the agent’s trajectory aligns with the reference path (Ilharco et al., 2019; Krantz et al., 2020; Raychaudhuri et al., 2021; 2025). Tasks involving classification or detection often use precision (how often predictions are correct), recall (how well all relevant targets are found), F1 score (combines and balances precision and recall) and receiver-operating characteristic curve or ROC curve (measures the tradeoff between sensitivity or true positive rate and specificity or false positive rate) (Chen & Mooney, 2011; Magnusson et al., 2009; Matuszek et al., 2012; Tellex et al., 2011b; Wellhausen et al., 2020) to evaluate prediction quality. In exploration tasks, coverage is used to measure the proportion of the environment observed (Chaplot et al., 2019). In manipulation tasks such as object rearrangement or tabletop operations, task success rate, measuring the percentage of trials where the agent achieves the desired manipulation (e.g., grasping, placing), is the dominant metric (Kalashnikov et al., 2018; Batra et al., 2020a; Lin et al., 2020; Zeng et al., 2020; Huang et al., 2023b; Gu et al., 2023). Some works also evaluate policy outputs using precision (Zeng et al., 2017), or employ pose error metrics to assess deviations between the final and target object positions (Liu et al., 2012).
@@ 6.2 Intrinsic evaluation
Accuracy. Map accuracy refers to how accurately the map captures geometric or semantic information when compared against the ground truth. However obtaining the ground truth map can be challenging in most cases. Some works (Georgakis et al., 2022a; Ramakrishnan et al., 2020) evaluate the 2D occupancy maps by using map accuracy to measure how much of that area matches with the ground-truth, and IoU to measure intersection over union between the built and the ground-truth map. While others (Cartillier et al., 2021; Georgakis et al., 2022b) evaluate the accuracy of the built 2D semantic map using semantic segmentation metrics such as pixel-wise labeling accuracy, pixel-based F1 score, IoU score and a contour-based average boundary F1 score. However, there is a lack of standardized metrics that can be used across various map types. For example, IoU is not directly applicable to topological maps. Moreover, while evaluating accuracy for semantic maps that store a fixed set of object categories is straightforward, it is tricky to evaluate semantic maps that store open-vocabulary features. As an alternative, OpenScene (Peng et al., 2023) performs semantic segmentation on their open-vocabulary map for a fixed set of categories and thereafter evaluate it using accuracy and IoU. ConceptGraphs (Gu et al., 2024), on the other hand, evaluate their open-vocabulary topological map by employing human evaluators on Amazon Mechanical Turk (AMT) to report the scene graph accuracy. OpenLex3D (Kassab et al., 2025) has been recently introduced to evaluate open-vocabulary 3D scene representations, by providing object labels at various granularity or precision. Similar evaluation techniques could benefit intrinsic evaluation in embodied setting as well. The challenge to evaluate modern-day open-vocabulary maps lies in collecting precise ground-truth maps for large real-world environments, which is labor-intensive or infeasible. This limits the use of direct comparisons in order to calculate map accuracy. Moreover, accuracy in dynamic or cluttered environments is also rarely assessed. Both these provide possible future area of research.
@@
Precision, Recall, F1, ROC Language grounding (2011; 2012; 2011b)